# UNIFY: Unified Index for Range Filtered Approximate Nearest Neighbors Search

### v1

@@ 4.3. Theoretical Analysis
Space Complexity. Following the original work of HNSW (Malkov and Yashunin, 2018), we use a 32-bit integer to store an edge in HSIG and analyze space complexity using 32-bit as a storage unit. In HSIG, each node has up to $MS$ edges (taking $MS$ storage units), a bitmap of size $MS$ (taking $\frac{MS}{32}$ units), and a skip list connection (taking one unit). For a dataset of $n$ objects, the total number of nodes in HSIG is $nL^{\prime}$, where $L^{\prime}$ represents the average number of levels in HSIG. This results in an expected space complexity of $O(nL^{\prime}(\frac{33}{32}MS+1))$. As discussed in (Malkov and Yashunin, 2018), the average number of levels in HNSW is a constant, and since HSIG uses the same strategy to determine the number of levels, $L^{\prime}$ is also a constant in HSIG. The values of $S$ and $M$ are determined experimentally and are generally small constants. Given that $L^{\prime}$, $S$, and $M$ are all considered small constants relative to the dataset size $n$, the space complexity can be simplified to $O(n)$.
Construction Complexity. Consider a dataset with $n$ objects divided into $S$ subsets, each containing $n_{1},n_{2},\dots,n_{S}$ objects. Inserting an object into HSIG requires three operations: (OP1) backbone graph insertion, (OP2) skip list insertion, and (OP3) edge masking. OP3 can be completed in constant time, while OP2 has an expected time complexity of $O(\log n)$ (Pugh, 1990). OP1 involves performing ANNS on the HNSW in each segment, totaling $S$ iterations of ANNS. Since ANNS on an HNSW with $x$ objects takes $O(\log x)$ time, the total cost of OP1 is $\sum_{i=1}^{S}\log(n_{i})$, which is bounded by $\sum_{i=1}^{S}\log n=S\log n$. Therefore, OP1 is an $O(S\log n)$ operation. Neglecting constants, the expected insertion complexity of adding an object is $O(\log n)$, and constructing an HSIG for $n$ objects scales as $O(n\log n)$.
Search Complexity. The complexity scaling of a single search can be strictly analyzed under the assumption that HSIG is able to exactly guarantee inclusivity with respect to HNSW. For an HSIG with $n$ objects, consider a query range covering $Y$ objects and a small constant $k$ that is negligible compared to $n$. The search complexities of the three strategies in HSIG are as follows:
A. Pre-Filtering. Searching the skip list has an expected complexity of $O(\log n)$ (Pugh, 1990), and computing vector distances within the query range takes $O(Y)$ time. Thus, the total time complexity is $O(Y+\log n)$. Due to the range-aware search in HSIG, Pre-Filtering is only employed when $Y$ is smaller than a constant $\tau_{A}$, ensuring that Pre-Filtering typically operates with an $O(\log n)$ complexity.
@@
C. Hybrid Filtering. For RF-ANNS, the hybrid-filtering strategy identifies segments covering the query range and performs ANNS on the HNSW of these segments. Assuming there are $n^{\prime}$ objects in the intersected segments, searching the HNSW takes $O(\log n^{\prime})$ time. Since $n^{\prime}\leq n$, the final expected time complexity is $O(\log n)$. Although the theoretical complexity is derived under the exact inclusivity assumption, experimental results (Section 5.8) confirm the method’s logarithmic scaling with increasing data size, validating its efficiency and scalability for large datasets.
5. EXPERIMENT
5.1. Experimental Setup
Datasets. We use six real-world datasets of varying sizes and dimensions in experiments. The Paper (Wang et al., 2023) and WIT-Image (Zuo et al., 2024) datasets contain both feature vectors and attribute values. The Paper includes publication, topic, and affiliation attributes, which we convert from categorical to numerical, while WIT-Image uses image size as its attribute. For the remaining datasets, which only contain feature vectors, we generate numerical attributes for each object using an attribute generation method similar to that described in (Wang et al., 2021b), augmenting each vector with a randomly generated attribute value between 0 and 10,000. The characteristics of the datasets are detailed in Table 2. For each query, we generate a query range uniformly between 0.1% and 100%.
@@ 5.1. Experimental Setup
The code for most methods is publicly accessible online. For methods without available code, we implemented them based on their descriptions in the respective papers. Since NHQ initially only supported attribute matching, we modified its fusion distance computation to account for the absolute differences between attribute values and applied the post-filtering strategy for RF-ANNS. We use the Euclidean distance function to measure vector distances.
@@
Parameter Settings. There are three crucial parameters in HSIG: $S$, $M$, and $efCons$, representing the number of segments, the maximum number of edge connections, and the number of candidate neighbors during index construction, respectively. We use grid search to find the optimal values, setting $S$, $M$, and $efCons$ to 8, 16, and 500, respectively. The parameters $m$ and $ef$ relate to the search process, where $m$ denotes the total number of neighbors visited per object, and $ef$ represents the number of candidates searched during a query. We vary the values of $m$ and $ef$ to generate the recall/QPS curves. We also use grid search to set parameters of baselines.
Implementation Settings. We implement the core parts of the HSIG construction and search algorithms based on hnswlib (Malkov and Yashunin, 2018). The code is written in C++ and compiled with GCC 10.3.1 using the ”-O3” optimization flag. We provide a Python interface for the core indexing library and conduct experiments using Python 3.8.17.
Environment. The scalability experiments are conducted on Alibaba Cloud Linux 3.2104 LTS with 40 physical cores and 512GB memory. The remaining experiments are conducted on a Linux server with the following hardware specifications: an Intel(R) Xeon(R) CPU E5-2609 v3 at 1.90GHz with 6 physical cores, 16GB memory, and the Ubuntu 18.04.5 operating system.
@@ 5.2. Overall Performance
We evaluate the query performance of HSIG and its competitors with $k$ values of 10 and 100. Based on the preprocessing method in Section 4.2, we set the default values of $\tau_{A}$ and $\tau_{B}$ to 1% and 50% of the dataset size, respectively. Figure 5 displays the query performance, and Table 2 shows the index sizes and build times. Although HSIG does not excel in index size or build time due to its multi-segment structure and extensive edge connections, it is comparable to SeRF, the state-of-the-art PG-based solution for RF-ANNS. PQ-based methods like ADBV and Milvus are space-efficient but struggle to deliver competitive query accuracy and efficiency compared to PG-based methods. As shown in Figure 5, HSIG consistently outperforms the baselines across all datasets regarding the QPS vs. recall trade-off. For example, with $k=10$ and a recall of around 0.9, HSIG achieves a QPS two orders of magnitude higher than ADBV on the GloVe and GIST1M datasets, one order of magnitude higher than NHQ on the SIFT1M dataset, and outperforms SeRF by up to 2.29 times across all datasets. These results highlight HSIG’s effectiveness due to its unified graph structure and range-aware strategy selection. Additionally, HSIG allows HNSW reconstruction with varying edge degrees using the parameter $m$, whereas SeRF reconstructs HNSW with a fixed edge degree, limiting its query performance. Finally, HSIG outperforms NGT, Vearch, and NHQ, as they employ the post-filtering strategy and perform poorly on small query ranges.
@@ 5.3. Effect of Range-aware Search Strategy Selection
In this section, we evaluate HSIG’s performance across small, medium, and large query ranges. The methods compared are as follows: (1) HSIG-pre executes searches on HSIG using the pre-filtering strategy. (2) PQ-pre performs attribute filtering first, followed by PQ-based vector retrieval. (3) Btree-pre uses a B-tree for attribute filtering, followed by brute-force vector retrieval. (4) HSIG-post applies HSIG with the post-filtering strategy. (5) HNSW-post builds HNSW over the entire dataset and performs RF-ANNS using the post-filtering strategy. (6) HSIG-hybrid employs HSIG for RF-ANNS queries using the hybrid filtering strategy. (7) SeRF is a state-of-the-art RF-ANNS solution that uses a hybrid filtering strategy. (8) HSIG-range-aware conducts RF-ANNS using range-aware search strategy selection. (9) Dedicated builds specialized indexes for each strategy, with Btree-pre, HNSW-post, and SeRF used for pre-, post-, and hybrid filtering, respectively, and selects the best strategy based on the query range.
The results are shown in Figure 6. Some methods have missing QPS values for specific query ranges, indicating they could not meet the recall threshold. Notably, HSIG-pre, HSIG-hybrid, and HSIG-post outperform their competitors in small, medium, and large ranges, respectively. For example, HSIG-pre outperforms Btree-pre by 20.3% at a recall of 0.9 in small query ranges. HSIG-hybrid exceeds SeRF by 1.21 times in medium query ranges at a recall of 0.95. HSIG-post outperforms HNSW-post by 37.5% at a recall of 0.99 in large query ranges. Additionally, HSIG-pre outperforms HSIG-hybrid and HSIG-post in small ranges, HSIG-post surpasses HSIG-hybrid and HSIG-pre in large ranges, and HSIG-hybrid performs best among the three strategies in medium ranges, validating the effectiveness of the proposed heuristics for range-aware strategy selection. Finally, HSIG-range-aware consistently outperforms Dedicated by up to 1.1 times across all query ranges, thanks to its unified PG-based index and range-aware strategy selection.
@@ 5.4. Validation of Inclusivity of HSIG
To further evaluate the impact of inclusivity on query performance, we compare HSIG at varying levels of inclusiveness (30%, 40%, 60%, and 80%) against two competitive methods that guarantee exact inclusivity. The first method, Optimal HNSW, builds an HNSW in real time for objects within each query range. The second method, MS-HNSW, pre-builds HNSWs for each segment. During the search, MS-HNSW identifies the segments intersecting with the query range, retrieves vectors from the corresponding HNSWs, and combines the intermediate results to obtain the final results. The results are shown in Figure 8. While Optimal HNSW offers the best performance, building indexes in real time for every query is time-consuming and impractical. As HSIG’s inclusiveness increases, its query performance improves, approaching that of Optimal HNSW. Additionally, HSIG outperforms MS-HNSW, which requires more distance computations. These results demonstrate the effectiveness of HSIG’s inclusivity, showing that higher inclusiveness leads to better query performance.
@@ 5.5. Validation of Incremental Insertion
In this section, we evaluate HSIG’s incremental insertion capability. We first build HSIG with 200,000 objects from the SIFT1M dataset, followed by four rounds of incremental insertion, each adding 200,000 objects. We compare the index build time and query performance against the state-of-the-art method SeRF. As shown in Figure 9, HSIG’s build time remains stable with each insertion since the number of inserted objects is consistent, whereas SeRF’s build time increases linearly due to its lack of incremental update support. Moreover, HSIG outperforms SeRF in query efficiency at a recall of 0.95. These results highlight HSIG’s effectiveness in supporting incremental insertions, showing it is suitable for applications with continuously evolving data.
Figure 9. Performance of incremental insertion on SIFT1M.
@@ 5.7. Parameter Study
In this section, we analyze the sensitivity of key parameters involved in both index construction and search performance.
Impact of Index Construction Parameters. We analyze the sensitivity of three parameters in HSIG construction: $S$, $M$, and $efCons$. Figures 11 and 12 illustrate how these parameters affect query performance and index build times, respectively. Figure 11a shows the impact of varying $S$ on query performance. We observe a gradual performance improvement as $S$ increases from 2 to 8. However, further increasing $S$ results in a decline in performance because more segments require visiting more neighbors per object during the search. Figure 11b and Figure 11c illustrate the effects of varying $M$ and $efCons$ on query performance, respectively. A lower $M$ degrades the quality of the graph structure, while a higher $M$ increases the number of objects traversed during the search. Similarly, a lower $efCons$ results in insufficient candidate neighbors, whereas a higher $efCons$ includes more irrelevant candidates, both hindering query performance. Therefore, setting these parameters based on the dataset and workload is crucial for balancing query efficiency and accuracy. As shown in Figure 12, index build time increases with higher values of $S$, $M$, and $efCons$ due to the increased distance computations required during construction. Based on a comprehensive evaluation of query performance and index construction time, we select $S=8$, $M=16$, and $efCons=500$ as the default parameter settings.
Figure 11. Performance of index parameters on SIFT1M.
@@
Impact of $k$ values. Figure 14 shows the impact of different $k$ values on the hybrid filtering performance of HSIG. HSIG maintains strong query efficiency and accuracy across various $k$ values. However, as $k$ increases from 10 to 100, performance gradually declines due to the increased number of candidates that need to be filtered during the search.
@@ 5.9. Discussions
Range-aware strategy selection. As mentioned in Section 4.2, our range-aware strategy selection method is based on historical data statistics, not query patterns. This approach may face challenges if the data distribution of the base dataset changes significantly over time. To address this, we propose an adaptive method that can detect changes in data distribution. If the change exceeds the user-defined threshold, it resamples objects from the updated base dataset to recalibrate $\tau_{A}$ and $\tau_{B}$, thereby adjusting the range-aware search strategy selection.
RF-ANNS with Multiple Attributes. Existing PG-based indexes struggle to support RF-ANNS queries with multiple attributes, as incorporating multiple attributes into a graph is challenging. However, with two enhancements, HSIG can be extended to handle such queries. While we focus on the case with two attributes in this discussion, the approach can be easily adapted for more attributes. (1) Multiple single-attribute indexes: Build a separate HSIG for each attribute. For a conjunctive query with a query vector $q$, retrieve objects that satisfy $r_{1}(A_{1})$ AND $r_{2}(A_{2})$, where $r_{1}(A_{1})$ and $r_{2}(A_{2})$ are the attribute ranges. The goal is to return the top-$k$ ANN of $q$ among objects satisfying both ranges. Specifically, we modify Line 15 of Algorithm 7 to ”if $v$ satisfies $r_{1}(A_{1})$ AND $r_{2}(A_{2})$, then push $v$ to $ann$”. For disjunctive queries ($r_{1}(A_{1})$ OR $r_{2}(A_{2})$), we use separate indexes for $A_{1}$ and $A_{2}$ and merge the results. (2) Single index for multiple attributes: Create a composite attribute $(A_{1},A_{2})$ and apply the z-order method (Orenstein and Merrett, 1984) to map them into a one-dimensional attribute, enabling the construction of a single HSIG to handle RF-ANNS queries with multiple attributes. We plan to explore a dedicated algorithm for RF-ANNS queries with multiple attributes in future work.
@@
Figure 15. Impact of varying data size on SIFT dataset.

### v2

@@ 4.3. Theoretical Analysis
Space Complexity. Following HNSW (Malkov and Yashunin, 2018), we use a 32-bit integer to store each edge in HSIG and analyze space complexity using 32-bit as a storage unit. In HSIG, each node has up to $MS$ edges (taking $MS$ units), a bitmap of size $MS$ (taking $\frac{MS}{32}$ units), and a skip list connection (taking one unit). For a dataset of $n$ objects, HSIG contains $nL^{\prime}$ nodes, where $L^{\prime}$ is the average number of levels. This results in an expected space complexity of $O(nL^{\prime}(\frac{33}{32}MS+1))$. As discussed in (Malkov and Yashunin, 2018), the average number of levels in HNSW is a constant, and since HSIG uses the same strategy to determine the number of levels, $L^{\prime}$ is also a constant in HSIG. $S$ and $M$ are experimentally determined and are generally small constants. Given that $L^{\prime}$, $S$, and $M$ are all considered small constants relative to $n$, the space complexity can be simplified to $O(n)$.
Construction Complexity. Consider a dataset with $n$ objects divided into $S$ subsets, each containing $n_{1},n_{2},\dots,n_{S}$ objects. Inserting an object into HSIG requires three operations: (OP1) backbone graph insertion, (OP2) skip list insertion, and (OP3) edge masking. OP3 runs in constant time, while OP2 has an expected complexity of $O(\log n)$ (Pugh, 1990). OP1 performs ANNS on the HNSW of each segment, requiring $S$ ANNS iterations. Since ANNS on an HNSW with $x$ objects takes $O(\log x)$ time, OP1 costs $\sum_{i=1}^{S}\log(n_{i})$ bounded by $S\log n$, making OP1 an $O(S\log n)$ operation. Neglecting constants, the expected complexity of inserting an object is $O(\log n)$, and constructing an HSIG for $n$ objects scales as $O(n\log n)$.
Search Complexity. The complexity scaling of a single search can be strictly analyzed under the assumption that HSIG exactly guarantees inclusivity with respect to HNSW. For an HSIG with $n$ objects, consider a query range covering $Y$ objects and a small constant $k$ negligible compared to $n$. The search complexities of the three strategies in HSIG are as follows:
A. Pre-Filtering. Searching the skip list has an expected complexity of $O(\log n)$ (Pugh, 1990), and computing vector distances within the query range takes $O(Y)$ time. Thus, the total complexity is $O(Y+\log n)$. Due to range-aware search in HSIG, Pre-Filtering is used only when $Y$ is smaller than a constant $\tau_{A}$, ensuring it typically operates with an $O(\log n)$ complexity.
@@
C. Hybrid Filtering. For RF-ANNS, hybrid filtering selects segments covering the query range and performs ANNS on their HNSWs. Let $n^{\prime}$ be the number of objects in the intersected segments, searching the HNSW takes $O(\log n^{\prime})$ time. Since $n^{\prime}\leq n$, the expected time complexity is $O(\log n)$. Although the theoretical complexity is derived under the exact inclusivity assumption, experimental results (Section 5.8) confirm the method’s logarithmic scaling with data size, validating its efficiency and scalability.
@@
5. EXPERIMENT
5.1. Experimental Setup
Datasets. We use six real-world datasets of varying sizes and dimensions. The Paper (Wang et al., 2023) and WIT-Image (Zuo et al., 2024) datasets include both feature vectors and attributes. The Paper contains publication, topic, and affiliation attributes, which we convert from categorical to numerical, while WIT-Image uses image size as its attribute. For the remaining datasets, which contain only feature vectors, we generate numerical attributes using a method similar to (Wang et al., 2021b), assigning each vector a random value between 0 and 10,000. The dataset characteristics are detailed in Table 1. For each query, we generate a query range uniformly between 0.1% and 100%.
@@ 5.1. Experimental Setup
We use the Euclidean distance function to measure vector distances.
@@
Parameter Settings. The parameters $S$, $M$, and $efCons$ represent the number of segments, maximum edge connections, and candidate neighbors during index construction, respectively. We use grid search to determine their optimal values, setting $S$, $M$, and $efCons$ to 8, 16, and 500, respectively. The parameters $m$ and $ef$ relate to the search process, where $m$ is the total number of neighbors visited per object, and $ef$ is the number of candidates searched during a query. We vary $m$ and $ef$ to generate recall/QPS curves and apply grid search to set baseline parameters.
Implementation Settings. We implement HSIG construction and search algorithms based on hnswlib (Malkov and Yashunin, 2018). The code is written in C++ and compiled with GCC 10.3.1 using the ”-O3” optimization flag. A Python interface is provided for the indexing library, and experiments are conducted using Python 3.8.17.
Environment. Scalability experiments are conducted on Alibaba Cloud Linux 3.2104 LTS with 40 cores and 512GB memory. Other experiments are conducted on a Linux server with an Intel(R) Xeon(R) E5-2609 v3 (1.90GHz, 6 cores), 16GB memory, and Ubuntu 18.04.5.
@@ 5.2. Overall Performance
We evaluate the query performance of HSIG and its competitors with $k$ values of 10 and 100. Based on the preprocessing method in Section 4.2, we set $\tau_{A}$ and $\tau_{B}$ to 1% and 50% of the dataset size. Figure 5 shows query performance, while Table 2 presents index sizes and build times. Although HSIG does not excel in index size or build time due to its multi-segment structure and extensive edge connections, it is comparable to SeRF, the state-of-the-art PG-based solution for RF-ANNS. PQ-based methods like ADBV and Milvus are space-efficient but struggle with query accuracy and efficiency compared to PG-based methods. As shown in Figure 5, HSIG consistently outperforms baselines across all datasets regarding the QPS vs. recall trade-off. For example, with $k=10$ and a recall of around 0.9, HSIG achieves a QPS two orders of magnitude higher than ADBV on the GloVe and GIST1M datasets, one order of magnitude higher than NHQ on the SIFT1M dataset, and outperforms SeRF by up to 2.29 times across all datasets. These results highlight HSIG’s effectiveness, benefiting from its unified graph structure and range-aware strategy selection. Additionally, HSIG allows HNSW reconstruction with varying edge degrees using the parameter $m$, whereas SeRF reconstructs HNSW with a fixed edge degree, limiting its query performance. Finally, HSIG outperforms NGT, Vearch, and NHQ, as they employ the post-filtering strategy and perform poorly on small query ranges.
@@ 5.3. Effect of Range-aware Search Strategy Selection
We evaluate HSIG’s performance across small, medium, and large query ranges. The methods compared are as follows: (1) HSIG-pre uses HSIG with the pre-filtering strategy. (2) PQ-pre performs attribute filtering first, followed by PQ-based vector retrieval. (3) Btree-pre uses a B-tree for attribute filtering, followed by brute-force vector retrieval. (4) HSIG-post applies HSIG with the post-filtering strategy. (5) HNSW-post builds HNSW over the entire dataset and applies the post-filtering strategy. (6) HSIG-hybrid employs HSIG for RF-ANNS queries using the hybrid filtering strategy. (7) SeRF is a state-of-the-art RF-ANNS solution that uses a hybrid filtering. (8) HSIG-range-aware uses HSIG with range-aware search strategy selection. (9) Dedicated builds specialized indexes for each strategy, with Btree-pre, HNSW-post, and SeRF used for pre-, post-, and hybrid filtering, respectively, and selects the best strategy based on query range.
The results are shown in Figure 6. Some methods have missing QPS values for specific query ranges, indicating they could not meet the recall threshold. HSIG-pre, HSIG-hybrid, and HSIG-post outperform competitors in small, medium, and large ranges, respectively. For example, HSIG-pre outperforms Btree-pre by 20.3% at a recall of 0.9 in small query ranges. HSIG-hybrid exceeds SeRF by 1.21 times in medium query ranges at a recall of 0.95. HSIG-post surpasses HNSW-post by 37.5% at a recall of 0.99 in large query ranges. Additionally, HSIG-pre outperforms HSIG-hybrid and HSIG-post in small ranges, HSIG-post surpasses HSIG-hybrid and HSIG-pre in large ranges, and HSIG-hybrid performs best among the three strategies in medium ranges, validating our range-aware heuristics. Finally, HSIG-range-aware consistently outperforms Dedicated by up to 1.1 times across all query ranges, thanks to its unified PG-based index and range-aware strategy selection.
@@ 5.4. Validation of Inclusivity of HSIG
To further evaluate the impact of inclusivity on query performance, we compare HSIG at varying levels of inclusiveness (30%, 40%, 60%, and 80%) against two competitive methods that guarantee exact inclusivity. The first method, Optimal HNSW, builds an HNSW in real time for objects within each query range. The second method, MS-HNSW, pre-builds HNSWs for each segment. During the search, MS-HNSW identifies the segments intersecting with the query range, retrieves vectors from the corresponding HNSWs, and combines the intermediate results to obtain the final results. The results are shown in Figure 8. While Optimal HNSW offers the best performance, building indexes in real time for every query is time-consuming and impractical. As HSIG’s inclusiveness increases, its query performance improves, approaching that of Optimal HNSW. Additionally, HSIG consistently outperforms MS-HNSW, which requires more distance computations. These results demonstrate the effectiveness of HSIG’s inclusivity, showing that higher inclusiveness leads to better query performance.
@@ 5.5. Validation of Incremental Insertion
In this section, we evaluate HSIG’s incremental insertion capability. We first build HSIG with 200,000 objects from the SIFT1M dataset, followed by four rounds of incremental insertion, each adding 200,000 objects. We compare the index build time and query performance against the state-of-the-art method SeRF. As shown in Figure 9, HSIG’s build time remains stable with each insertion since the number of inserted objects is consistent, whereas SeRF’s build time increases linearly due to its lack of incremental update support. Moreover, HSIG outperforms SeRF in query efficiency at a recall of 0.95. These results highlight HSIG’s effectiveness in supporting incremental insertions, showing that it is suitable for applications with continuously evolving data.
@@ 5.7. Parameter Study
Impact of Index Construction Parameters. We analyze the sensitivity of three parameters in HSIG construction: $S$, $M$, and $efCons$. Figures 11 and 12 show how parameters affect query performance and index build times, respectively. Figure 11a shows the impact of varying $S$ on query performance, with a performance increase from 2 to 8, followed by a decline as $S$ grows further due to increased neighbor visits. Figure 11b and Figure 11c illustrate the effects of varying $M$ and $efCons$ on query performance. Lower $M$ degrades graph quality, while higher $M$ increases the number of objects traversed during the search. Similarly, lower $efCons$ leads to insufficient candidates, whereas higher $efCons$ includes irrelevant candidates, both hindering query performance. Therefore, setting these parameters based on the dataset and workload is crucial for balancing query efficiency and accuracy. As shown in Figure 12, index build time rises with higher $S$, $M$, and $efCons$ due to increased distance computations. Based on a comprehensive evaluation of query performance and index construction time, we select $S=8$, $M=16$, and $efCons=500$ as default settings.
@@
Impact of $k$ values. Figure 14 shows the impact of different $k$ values on query performance of HSIG. HSIG maintains strong efficiency and accuracy across various $k$ values. However, as $k$ increases from 10 to 100, performance gradually declines due to the increased number of candidates that need to be filtered during the search.
@@ 5.9. Discussions
Range-aware strategy selection. As mentioned in Section 4.2, our range-aware strategy selection method is based on historical data statistics, not query patterns. This approach may face challenges if the data distribution of the base dataset changes significantly over time. To address this issue, we propose an adaptive method that can detect changes in data distribution. If the change exceeds the user-defined threshold, the method resamples objects from the updated base dataset to recalibrate $\tau_{A}$ and $\tau_{B}$, thereby adjusting the range-aware search strategy selection.
RF-ANNS with Multiple Attributes. Existing PG-based indexes struggle to support RF-ANNS queries with multiple attributes, as incorporating multiple attributes into a graph is challenging. However, with two enhancements, HSIG can be extended to handle such queries. While we use the case with two attributes in this discussion, the method can be adapted for more attributes. (1) Multiple single-attribute indexes: Build a separate HSIG for each attribute. For a conjunctive query with a query vector $q$, retrieve objects that satisfy $r_{1}(A_{1})$ AND $r_{2}(A_{2})$, where $r_{1}(A_{1})$ and $r_{2}(A_{2})$ are the attribute ranges. We aim to return the top-$k$ ANN of $q$ among objects satisfying both ranges. Specifically, we modify Line 15 of Algorithm 7 to ”if $v$ satisfies $r_{1}(A_{1})$ AND $r_{2}(A_{2})$, then push $v$ to $ann$”. For disjunctive queries ($r_{1}(A_{1})$ OR $r_{2}(A_{2})$), we use separate indexes for $A_{1}$ and $A_{2}$ and merge the results. (2) Single index for multiple attributes: Create a composite attribute $(A_{1},A_{2})$ and apply the z-order method (Orenstein and Merrett, 1984) to map them into a one-dimensional attribute, enabling the construction of a single HSIG to handle RF-ANNS queries with multiple attributes. We plan to explore a dedicated algorithm for RF-ANNS queries with multiple attributes in future work.